You are here

Beam Delivery News

Primary tabs

Title Date Excerpt Attached Image
Cyclotron workshop tomorrow, and…beam delivery? Now? 24 February 2015

There will be a workshop on the cyclotron tomorrow (February 25) in the Auditorium. This is geared towards operators but is open to anyone who may be interested.

Beam delivery update: Last update of 2014? 12 December 2014

The cyclotron is still delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A remains off while the next experiment is being set up.

Beam delivery update 11 December 2014

They cyclotron’s running fairly well, delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A is off at least overnight as we have no RIB users at ISAC.

Beam delivery update 10 December 2014

The cyclotron is delivering 9.5 µA to BL2A, ~95 µA to BL2C4 for isotope production, and beam as needed to BL1B for PIF. BL2A may go off for a few hours if RIB is not needed overnight, but otherwise we expect to continue like this for the next few days.

Beam delivery update 09 December 2014

Today’s planned beam development is still underway but is scheduled to wrap up by 20:30.

Beam delivery update: Short maintenance and beam development tomorrow 08 December 2014

The cyclotron is currently delivering ~110 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This will continue until ~07:00 tomorrow, at which point the cyclotron will go off for a short maintenance period followed by beam development.

Beam delivery update 05 December 2014

The BL1A current has been reduced to ~110 µA. BL2A is running at 9.8 µA while BL2C1 continues to take nA as needed for PIF. This is all expected to continue through the weekend.

Beam delivery update: Beam schedule changes 04 December 2014

The cyclotron’s delivering ~125 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This is expected to continue overnight and through tomorrow morning.

Beam delivery update 03 December 2014

The cyclotron is delivering ~125 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. At ISAC, 47K 46K is now being delivered to GRIFFIN on a discretionary basis. This is expected to continue for the next few days.

Beam delivery update: ISAC target plans 02 December 2014

The cyclotron is running normally, delivering ~125 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This is scheduled to continue overnight and though the coming days.

Beam delivery update: ISAC target failure and maintenance tomorrow 01 December 2014

The cyclotron is currently off due to an ISIS trip but it was running until about 21:20, delivering ~125 µA to BL1A and 9.8 µA to BL2A. It is scheduled to deliver to BL1A and BL2A overnight, with ~4 hours maintenance to start at 08:00 tomorrow.

Beam delivery update 26 November 2014

The cyclotron’s currently delivering ~125 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue at least overnight.

Proton therapy – update 26 November 2014

There will be only four treatments after all, one each on Friday, Saturday, Sunday, and Monday. Treatments will start at ~08:00 as usual and are expected to take ~2 hours.

Beam delivery update 25 November 2014

All three beamlines are now online, and the cyclotron is currently delivering ~115 µA to BL1A, 5 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight.

Beam delivery update: 4-hour maintenance tomorrow 24 November 2014

The cyclotron’s running normally, delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. Beam will go off at 08:30 tomorrow morning for ~4 hours of maintenance (and potentially more at ISAC to accommodate Target Hall work).

Beam delivery update: Weekend edition 21 November 2014

PIF wrapped up this afternoon and the cyclotron is delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue through the weekend.

Beam delivery update 20 November 2014

The cyclotron is delivering ~110 µA to BL1A and nA as needed to BL2C1 for PIF. BL2A has been off since about 07:00 this morning for target hall work.

Beam delivery update 19 November 2014

The cyclotron is currently delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4; this is expected to continue until about 07:00, at which point BL2A will go off for ISAC target hall work.

Beam delivery update 18 November 2014

The cyclotron is currently delivering ~100 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF; that is expected to continue overnight, with PIF resuming in the morning.

Beam delivery update 17 November 2014

The cyclotron is running and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. There is no cyclotron maintenance planned for tomorrow, but BL2A will go off around 07:00 for Target Hall work.

Beam delivery update: Weekend update 14 November 2014

The cyclotron is running with BL2A and BL2C4 both online. It’s not clear how long it will take for the BL1A vacuum to recover to where it can also be brought online; the current estimate is some time tomorrow (Saturday).

Beam delivery update: Maintenance underway 13 November 2014

The cyclotron is off for maintenance and will remain off overnight. It’s not clear at this point when either the cyclotron or BL1A will be ready for operation.

Beam delivery update: Cyclotron defrost tomorrow 12 November 2014

The cyclotron’s currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~40 µA to BL2C4, while at ISAC 8Li is still being delivered to βNMR/NQR. There are three shifts of maintenance scheduled starting tomorrow to accommodate a defrost of the cyclotron cryopanels and cryogenics system.

Beam delivery update: Lab closed tomorrow; long maintenance Thursday 10 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~50 µA to BL2C4. 8Li is being delivered to βNMR/NQR at ISAC for materials science experiments. This is expected to continue with only minor changes until Thursday morning.

Beam delivery update: Weekend edition 07 November 2014

The cyclotron is delivering ~110 µA to BL1A and 50 µA to BL2A. We’re scheduled to run with high currents in all three beamlines over the weekend and into next week.

Beam delivery update 06 November 2014

The BL2A current was reduced to 50 µA earlier today and BL2C4 is currently running at 80 µA, but BL1A is still running at ~110 µA and 8Li is still being delivered to βNMR/NQR at ISAC.

Beam delivery update 05 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue through the weekend.

Beam delivery update: ISAC maintenance tomorrow 04 November 2014

The cyclotron’s currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue until ~08:30 tomorrow morning at which time BL2A will be defined off for a target module move.

Beam delivery update: No maintenance tomorrow; ISAC maintenance Wednesday 03 November 2014

The cyclotron is currently delivering ~110 µA to BL1A, 60 µA to BL2A, and ~90 µA to BL2C4 (limited by the resistivity of the STF cooling water). This is expected to continue until Wednesday morning.

Beam delivery update 31 October 2014

The cyclotron is still delivering ~115 µA of beam to BL1A, 60 µA to BL2A, and ~95 µA to BL2C4. The BL2A current might still be adjusted, but otherwise we expect to run like this through the weekend.

Beam delivery update 30 October 2014

The cyclotron’s delivering ~115 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. There’s a chance that the BL2A current will be increased tomorrow as we switch to a new RIB experiment, but otherwise we expect to continue like this through the weekend.

Beam delivery update 29 October 2014

The cyclotron is currently delivering ~115 µA to BL1A, 50 µA to BL2A, and nA as needed to BL2C1 for PIF. This is expected to continue until ~20:00 at which point delivery to STF will resume.

Beam delivery update: Maintenance finished but startup delayed 28 October 2014

Today’s scheduled work has been completed and beam was injected late this afternoon; however, the cyclotron is off right now and it’s not clear when beam might actually be available.

Beam delivery update: Maintenance tomorrow 27 October 2014

The cyclotron is running normally and delivering ~115 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until the start of tomorrow's scheduled maintenance shift at 08:30.

Beam delivery update: Weekend edition 25 October 2014

BL2A was off for somewhat more than 15 minutes this morning to allow for access to the electrical services room, then for another hour early in the afternoon because of a fire alarm in the ISAC-I hall. BL2C was also off for about two and a half hours for access to the STF water package. The cyclotron seemed to behave well, at least, and is currently delivering ~115 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4.

Beam delivery update 23 October 2014

The cyclotron is continuing to deliver ~115 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4, while at ISAC 8Li delivery to βNMR and/or βNQR is ongoing. These are all expected to continue through the weekend.

Beam delivery update: Development finished; operation resuming 22 October 2014

The cyclotron is running normally and delivering ~115 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4.

Beam delivery update: Development finished; operation resuming 21 October 2014

Today’s scheduled beam development has finished and operation is resuming. The cyclotron is currently delivering ~105 µA to BL1A, 40 µA to BL2A (as the 2A current is still being stepped up), and ~95 µA to BL2C4.

Beam delivery update: Beam development tomorrow 20 October 2014

The cyclotron is currently delivering ~110 µA to BL1A and 50 µA to BL2A. This is expected to continue until the start of tomorrow’s planned beam development shift. Development is scheduled to run from 08:30 until roughly 20:00, at which time normal operation will be restored.

Beam delivery update: major power bump Saturday 19 October 2014

As many of you already know, the lab experienced a major power bump at about 09:45 yesterday. This had a big impact on almost every accelerator-related system on site: vacuum, cryogenics, RF, controls, cooling systems, building services, etc.

Beam delivery update 17 October 2014

The cyclotron is presently delivering ~110 µA to BL1A and 50 µA to BL2A; BL2C1 is set up for proton therapy tomorrow morning. High-current delivery to BL1A and BL2A, interrupted each morning for proton therapy, is expected to continue through the weekend.

Beam delivery update: Proton therapy starting tomorrow 16 October 2014

The cyclotron is delivering ~110 µA to BL1A and 50 µA to BL2A. This is expected to continue until the start of proton therapy at 08:00 tomorrow morning. BL2C1 was set up for PT this afternoon.

Beam delivery update 15 October 2014

The cyclotron is now running pretty much normally, delivering ~110 µA to BL1A and 50 µA to BL2A. The STF group hasn’t yet decided if they’re going to request beam overnight or if BL2C will stay off until tomorrow's switchover to 2C1.

Further to last night's post… 14 October 2014

The main magnet is on, the RF has been tested, and inflector/deflector conditioning is underway. The present hope is to inject beam some time later today for full production tomorrow as originally scheduled.

Mini-shutdown update: Cyclotron startup delayed 13 October 2014

A power bump due to generator testing tripped off the vacuum pump on the space between the cyclotron tank seals on Friday, causing a partial vent of the tank. The vacuum is still recovering. It looks like Wednesday is the earliest at which beam might be injected as a result.

Mini-shutdown update 09 October 2014

We’re about half-way through the scheduled mini-shutdown. Injection and startup are scheduled for Tuesday with beam delivery to resume Wednesday.

Beam delivery update 01 October 2014

The cyclotron is delivering ~105 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue at least overnight and possibly through the weekend.

Beam delivery update: ISAC maintenance tomorrow 29 September 2014

The cyclotron is currently delivering ~105 µA to BL1A, 5 µA to BL2A (for end-of-target target scans), and ~95 µA to BL2C4. BL2A is expected to go off at some point this evening.

Beam delivery update 20 September 2014

BL1A is the only beamline running right now; it’s taking ~105 µA and will likely do so through the weekend. BL2A is off due to target issues while BL2C is off until tomorrow morning’s proton therapy.

Beam delivery update 18 September 2014

Beam is currently off due to RF problems. The plan is to deliver ~105 µA to BL1A, 50 µA to BL2A, and 95 µA to BL2C4 overnight before inserting the pepper pot for proton therapy at 08:00 tomorrow morning.

Beam delivery update 16 September 2014

The cyclotron is running and delivering ~105 µA to BL1A and ~95 µA to BL2C4. BL2A is just coming on following today’s shielding block move and stable-beam tuning to the ISAC yield station.

Beam delivery update: very little maintenance, if any, tomorrow [UPDATED] 15 September 2014

The cyclotron is running normally and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and 90–95 µA to BL2C4. We expect to run like this until 06:00 tomorrow. Cyclotron operation may be interrupted briefly tomorrow morning for maintenance but at this point there is nothing planned. 

Beam delivery update: Weekend edition 12 September 2014

The cyclotron is running normally and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. PIF is scheduled to wrap up this evening, at which point STF operation on BL2C4 will resume; otherwise, we expect to run like this through the weekend.

Beam delivery update 11 September 2014

The cyclotron’s delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight. 

Beam delivery update 10 September 2014

The cyclotron is currently running with ~20 µA in BL1A; the plan for the evening is to bring BL1A up to full current, then to bring BL2C4 back online. BL2A will remain off until tomorrow morning.

Beam delivery update: Recovery and the next ~48 hours 09 September 2014

After last night’s spontaneous start of a cryopanel defrost, the Cryogenics group came in to defrost and restart the tank cryogenics system overnight. It’s still not clear when beam might be injected – it could be anywhere between this evening and tomorrow morning. Operation over the next couple of days is still somewhat uncertain.

Beam delivery update [UPDATED] 08 September 2014

The cyclotron is still delivering nA as needed to BL1B, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production. PIF operation is scheduled to end tonight, but BL2A and 2C4 will continue to operate until the start of maintenance at 06:00 Wednesday.

Beam delivery update: Weekend edition [UPDATED] 05 September 2014

The cyclotron is delivering 9.8 µA to BL2A and ~95 µA to BL2C4, while BL1B has finished operation for the evening. PIF will resume tomorrow morning. Otherwise, we’ll continue like this through the weekend.

Beam delivery update 04 September 2014

The cyclotron is running much as it was last night, delivering nA to BL1B, 9.8 µA to BL2A, and ~95 µA to BL2C4. Operation is scheduled to carry on like this into the weekend.

Beam delivery update 03 September 2014

The cyclotron’s running normally, delivering nA as needed to BL1B for PIF, 9.8 µA to BL2A, and ~95 µA to BL2C4. We expect to run like this until 06:00 tomorrow morning, at which point BL2A will again be defined off to allow the experimenters to access the separator room to recover their target foil.

Beam delivery update: Maintenance tomorrow 02 September 2014

The cyclotron is currently delivering nA as needed to BL1B for PIF, 9.8 µA to BL2A for RIB production, and ~95 µA to BL2C4 for isotope production at STF. This is expected to continue until ~06:00 tomorrow, at which point BL2A will be defined off for access to the mass separator room and a target module move. BL1A and BL2C4 will continue to run until ~08:30.

Beam delivery update 31 August 2014

The cyclotron is delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4, while at ISAC 31Na is being delivered to Osaka experiment S1391. Where we have a problem is with the ISAC-I RFQ.

Beam delivery update (SPECIAL LONG WEEKEND EDITION) 29 August 2014

The cyclotron is running normally, delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. Other than the possibility of an STF target change at some point (which would see BL2C4 defined off and might cause brief disruptions on the other beamlines), this is expected to continue through the weekend.

Beam delivery update 28 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF. With the exception of an STF target change at some point, this is expected to continue through the upcoming long weekend.

Beam delivery update: beam development completed 26 August 2014

Today’s beam development was finished shortly before 20:00 and full operation is being restored. The plan is to deliver ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 as conditions allow.

Beam delivery update: Maintenance during beam development tomorrow 25 August 2014

The cyclotron’s running normally, delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4, while at ISAC 34Al is being delivered to TITAN for experiment S1240. These are both scheduled to continue until the start of tomorrow’s beam development shift at 08:30. There is no plan to interrupt operation for Wednesday’s “maintenance if required”.

Beam delivery update 21 August 2014

The cyclotron is running normally, delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~98 µA to BL2C4. At ISAC, the new UCtarget in TM4 saw its first protons last night. Both cyclotron operation and RIB delivery are scheduled to continue through the weekend.

Beam delivery update: Maintenance tomorrow 19 August 2014

The cyclotron is delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue until 08:30 tomorrow, at which point the cyclotron will go off for an ISIS backplate change. 

Beam delivery update: weekend edition 15 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. Ops is working to address a low water flow issue.

Beam delivery update: operation restored 13 August 2014

The cyclotron is back to full operation and delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. We’re scheduled to run like this until next week’s maintenance.

Cyclotron down due to water pump failure 13 August 2014

The RF system tripped off at about 06:15 this morning. On investigation, it was found that the north Cu LCW pump had failed, roughly halving the pressure in the LCW system. Ops has shut down other loads on the system, including the main magnet power supply, as a precaution against damage.

Beam delivery update 11 August 2014

At this moment, the cyclotron is running and delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4.

Beam delivery update: weekend edition 08 August 2014

The cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue through the weekend.

Beam delivery update 07 August 2014

The cyclotron is still delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF; this is still scheduled to continue through the weekend, as is 11Be delivery to TIGRESS for experiment S1429.

Beam delivery update 06 August 2014

The cyclotron is running normally, delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. Operation is scheduled to continue like this through the weekend.

Beam delivery update: Operation restored 05 August 2014

Operation was restored at about 17:45 and the cyclotron is currently delivering ~110 µA to BL1A, 50 µA to BL2A, and ~97 µA to BL2C4. This is expected to continue until roughly 08:00 tomorrow, at which point BL2C will be switched to PIF until sometime in the afternoon.

Beam delivery update: Maintenance today 05 August 2014

It was decided, despite the late start, to go ahead with a source filament and backplate change today. This is expected to take ~6 hours. Beam went off just before 13:00 and is expected back early this evening.

Beam delivery update: weekend edition 04 July 2014

The cyclotron is under vacuum, the repairs to the STF beamline window have been completed, and the inflector and deflector have been successfully tested at their operating voltages with the main magnet off. At this point, and assuming no unforeseen problems between now and then, it looks like beam will be injected some time tomorrow (Saturday).

Beam delivery…er, cyclotron repair update 03 July 2014

The removable section of the vertical injection beamline has been replaced and the cyclotron is being baked out overnight; the plan is to start cooling it back down at 05:00 so that the Cryogenics group can begin cooling the tank cryopanels later in the morning.

Beam delivery update: lid down; cyclotron pumping 02 July 2014

The cyclotron tank lid has been lowered and pumping is getting underway. At this point, it’s estimated that we might inject Saturday for delivery on Sunday.

Beam delivery update: lid up; cyclotron down 27 June 2014

The cyclotron has been vented and the lid raised to allow the electrical fault in the deflector negative plate to be investigated and fixed. The plan at this point is to lower the lid on Wednesday, July 2 and to resume cyclotron operation on or about Monday, July 7.

Beam delivery update: Cyclotron down tomorrow morning for major repairs 26 June 2014

The cyclotron is currently limping along, delivering 9.8 µA to BL2A and beam as needed to BL1B for PIF. This will continue until 06:00 tomorrow (Friday), at which point the cyclotron will be turned off to begin warming up the cryopanels in preparation for venting later in the morning.

Cyclotron down, time to recovery not clear (updated) 25 June 2014

The cyclotron has been down since about 11:30 this morning and will remain off at least overnight due to a problem with the deflector.

Beam delivery update 24 June 2014

We’re running, with 9.8 µA in BL2A and small amounts of beam as needed in BL1B for proton irradiation. BL2C4 is still offline due to a cooling water leak at the solid target facility (STF). Repairing that leak will require the cyclotron to be off for one to 1-1/2 days, likely next week.

Beam delivery update: Maintenance tomorrow 23 June 2014

The cyclotron’s currently delivering ~115 µA to BL1A and 9.8 µA to BL2A. BL2C4 has been off since shortly after 12:00 because of a leak in the STF cooling water system. There is maintenance planned for tomorrow.

Beam delivery update: weekend edition 20 June 2014

We’re currently delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. That is scheduled to continue until Tuesday’s if-needed maintenance.

Beam delivery update 19 June 2014

The cyclotron’s delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~90 µA to BL2C4. At ISAC, the effort to deliver 95Sr16+ to TIGRESS for experiment S1389 is underway.

Beam delivery update 18 June 2014

The cyclotron’s running relatively well and delivering ~115 µA to BL1A, 8 µA to BL2A, and ~95 µA to BL2C4. At some point the BL2A current may be pushed to 9.8 µA for RIB production at ISAC, but otherwise we expect to run like this for the time being.

Beam delivery update 17 June 2014

We’re currently running with ~110 µA in BL1A, ~2 µA in BL2A, and ~15 µA in BL2C4. The plan for overnight is to ramp up the BL2C4 current and, if RIB tuning at ISAC allows, to carry out a BL2A target scan in advance of increasing the 2A current tomorrow.

Beam delivery update: Long maintenance underway 16 June 2014

This week’s long maintenance period is underway. The Cryogenics/Vacuum group will work to restore normal vacuum in the cyclotron this evening and the cyclotron RF will be restored tomorrow. We still hop to resume operation and beam delivery late afternoon rather than tomorrow evening.

Beam delivery update: Long maintenance Monday 13 June 2014

The cyclotron is currently delivering ~115 µA to BL1A, 98 µA to BL2A, and nA as needed to BL2C1 for PIF. This will continue until ~08:30 tomorrow morning, at which point BL2A will be defined off for ~2 hours to allow access to the ISAC mass separator room. There is a 36-hour maintenance scheduled to begin at 08:30 Monday to accommodate a defrost of the cyclotron cryopanels.

Beam delivery update 10 June 2014

Tonight's plan is to resume operation with ~115 µA in BL1A, 70 µA in BL2A, and nA as needed in BL2C1 for PIF and to run like that for the time being.

Beam delivery update: Maintenance tomorrow 09 June 2014

The cyclotron is currently delivering ~118 µA to BL1A, 70 µA to BL2A, and nA as needed to BL2C1. Cyclotron maintenance will start at 10:00 tomorrow and is expected to take roughly two hours; ISAC, on the other hand, will go off-line at 08:30 for 6–8 hours.

Beam delivery update: Weekend edition 06 June 2014

The cyclotron is delivering ~96 µA to BL1A and 70 µA to BL2A. BL2C is set up for proton therapy over the weekend; beam will go off for about 90 minutes around 08:00–08:30 each morning from now through Monday.

Beam delivery update: Setup for PT tomorrow 04 June 2014

The cyclotron’s running normally, delivering ~90 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue until about 16:00 tomorrow afternoon, at which point we’ll switch from isotope production at STF to setup for proton therapy on BL2C1.

Beam delivery update 03 June 2014

Cyclotron beam development has been underway since 08:30. High-current operation will resume late this evening and continue until we set up for proton therapy Thursday afternoon.

Beam delivery update: Beam development Tuesday 02 June 2014

The cyclotron is delivering ~90 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue until the start of cyclotron beam development at 08:30 tomorrow.

Beam delivery update: weekend edition 30 May 2014

The cyclotron’s currently delivering ~90 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. At ISAC, 38K7+ is being delivered to DRAGON for experiment S1442. We should continue like this until the start of beam development on Tuesday.

Beam delivery update 28 May 2014

The cyclotron is currently delivering ~92 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4; this is expected to continue overnight and through tomorrow.

Beam delivery update: Maintenance tomorrow 26 May 2014

The cyclotron ran relatively smoothly over the weekend and is still delivering ~93 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. Maintenance is scheduled to start at ~08:30 tomorrow (Tuesday) morning and to take roughly eight hours.

Beam delivery update: Weekend edition 23 May 2014

The cyclotron is delivering ~95 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. At ISAC, 37K is being delivered to TRINAT for experiment S1188 while stable-beam setup of DRAGON experiment S1442 is carried out in parallel. These should all continue over the weekend.

Beam delivery update 22 May 2014

The cyclotron is delivering ~95 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. At ISAC,37K is being delivered to TRINAT for experiment S1188. Meanwhile, pilot-beam setup for experiment S1442 is nearly complete.

Beam delivery update 21 May 2014

The cyclotron is running relatively well, and we’re currently delivering ~90 µA to BL1A, 55 µA to BL2A, and ~95 µA to BL2C4. The BL2A current will be increased overnight as the ISAC target vacuum allows.

Beam delivery update: Long weekend edition 16 May 2014

The cyclotron is delivering ~95 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. BL2A will have to be defined off to allow access to the ISAC mass separator room at some point tomorrow, but otherwise delivery should continue like this until Tuesday morning.

Beam delivery update 15 May 2014

We’re currently running with ~95 µA in BL1A, 50 µA in BL2A, and ~95 µA in BL2C4. BL2C4 will be defined off at some point tomorrow for an STF target change. Otherwise, we should run like this until Tuesday morning.

Beam delivery update: Operation resuming 14 May 2014

Today’s maintenance is complete and operation is slowly being restored. We should be back at full operation, with high currents in each of BL1A, BL2A, and BL2C4, within the next hour or so.

Beam delivery update: Maintenance tomorrow 13 May 2014

The cyclotron’s still plugging along, delivering ~80 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This will continue until the start of maintenance at 08:30 tomorrow.

Beam delivery update: Full maintenance this week 12 May 2014

The cyclotron’s running relatively smoothly, delivering ~80 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. At ISAC, the RF buncher in the ISAC s-bend section tripped off at about 15:20 this afternoon and won't stay on. RF group is investigating. 

Beam delivery update 09 May 2014

PIF finished up early this afternoon and the cyclotron is now delivering ~80 µA to BL1A, 50 µA to BL2A, and ~92 µA to BL2C4. This is scheduled to continue until next Wednesday’s maintenance-if-required shift.

Beam delivery update 08 May 2014

The cyclotron is currently delivering ~80 µA to BL1A, 50 µA to BL2A, and nA as needed to BL2C1 for PIF. This is expected to continue overnight.

Beam delivery update 07 May 2014

We’re currently delivering ~80 µA to BL1A, 50 µA to BL2A, and ~75 µA to BL2C4; this is expected to continue overnight, with additional tuning likely tomorrow.

Beam delivery update 06 May 2014

Today’s scheduled maintenance is finished and the cyclotron is being started back up. The plan for the evening is to bring BL1A up to its full operating current of 100+ µA and to resume operation with BL2A and 2C4 overnight.

Beam delivery update: Maintenance tomorrow 05 May 2014

After a day’s worth of beam development, the cyclotron is now delivering 50 µA of protons to BL2A. Beam is expected to go off at ~08:30 tomorrow (Tuesday) for a scheduled ISIS backplate change.

Beam delivery update: Weekend edition [UPDATE] 03 May 2014

The Vancouver Marathon will take place tomorrow (Sunday) morning. There are a number of road closures that may affect access to the lab as a result.

Beam delivery update: Weekend edition 02 May 2014

The cyclotron is running and delivering 30 µA to BL2A and ~95 µA to BL2C4. BL2A will go off for 2–3 hours at ~08:30 tomorrow to allow access to the ISAC mass separator room for laser ion source setup

Beam delivery update 01 May 2014

The cyclotron is running and delivering 30 µA to BL2A, ~95 µA to BL2C4, and beam as needed to BL1B for PIF users. This is expected to continue until (at least) Saturday morning.

Beam delivery update: Proton therapy starts tomorrow 24 April 2014

BL1A is currently running with a trickle of beam (~3 µA) while 30 µA is being delivered to BL2A for ISAC RIB production. At ISAC, 8Li is being delivered to bNMR/NQR for a series of materials science experiments. 

Beam delivery update (including schedule changes) 23 April 2014

The cyclotron is running and delivering ~15 µA to BL1A, 30 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until 08:30 tomorrow morning, at which point proton currents will be reduced to <1 µA for spill tests in BL1A for about an hour. At ISAC, setup to bNMR/NQR is well underway and 15N3+ delivery to DRAGON for tests of new low-intensity diagnostics will continue until tomorrow evening.

Beam delivery update: Long weekend edition 17 April 2014

The cyclotron is running and delivering beam to all three high-current beamlines: ~20 µA to BL1A, 10 µA to BL2A2, and ~95 µA to BL2C4. This is expected to continue until some time Monday. A leak has been discovered in ISAC’s east target station (ITE); repairing the leak is expected to take three days shifts to be scheduled before the week of May 12.

Beam delivery update 16 April 2014

We’re about half-way through the Schedule 126 startup period. The cyclotron’s delivering ~25 µA to BL1A and ~98 µA to BL2C4. The BL2A vacuum has been sorted out and proton beam is expected in 2A at some point overnight. 

Schedule 126 update 15 April 2014

There have been a number of minor changes (of the housekeeping variety) made to Schedule 126 over the last two days…

Beam delivery update 14 April 2014

The 36Ar beam to TIGRESS was lost about three hours ago; the hope is to resume delivery before too long and to carry on through the week. The cyclotron’s currently delivering ~25 µA to BL1A and ~80 µA to BL2C4; BL2A will remain off overnight.

Beam delivery update: First beam of 2014 injected 11 April 2014

Beam was injected into the cyclotron a couple of hours ago. At ISAC, 36Ar7+, stripped to 13+ in HEBT, is being tuned to TIGRESS for experiment S1232.

Beam delivery update 10 April 2014

ISAC beam tuning is limited overnight to the RFQ and the first couple of DTL tanks as DTL Buncher 2 tripped off late this afternoon and could not be restarted. ISIS beamline tuning is underway in preparation for injection into the cyclotron.

Beam delivery update 09 April 2014

We’re delivering 4He at 25 MeV (6.25 MeV/u) to TIGRESS for detector calibrations, while on the cyclotron side of things end-of-shutdown activities are still being carried out.

Beam delivery update 08 April 2014

4He was delivered to TIGRESS for detector calibrations all day, and preparation for cyclotron startup is ongoing.

Beam delivery update 07 April 2014

The first beam of the year, 7.5 MeV/u 4He from the OLIS multicharge source, was delivered to TIGRESS for detector calibrations this afternoon. Shutdown work is continuing elsewhere.

Beam delivery update – it's that time of year again… 04 April 2014

Stable-beam operation at ISAC is scheduled to begin next week with delivery to TIGRESS. Tuning will continue over the weekend.

ISAC Electrical Panel P-373 Shutdown 16 January 2014

As part of the GRIFFIN installation, ISAC electrical panel P-373 needs to be shutdown for 4 hours at some point from January 28th to the 30th (during "core" working hours)…

Beam delivery update: Last update of 2013 18 December 2013

The cyclotron’s currently delivering 9.8 µA to BL2A, and beam as needed to BL1B and BL2C1 for PIF. This is scheduled to continue until Saturday morning, at which point things will be shut down for the year.

2013-12-16: Last week of operation in 2013 16 December 2013

This is the final week of accelerator operation and beam delivery in Schedule 125. The cyclotron is delivering 9.8 µA to BL2A and nA to BL2C1 as needed. RIB production and PIF (on both BL1B and 2C1) will continue until Saturday morning.

Beam delivery update: Last (operating) weekend of 2013 13 December 2013

The cyclotron is continuing to deliver 9.8 µA to BL2A and protons as needed to BL1B. This is expected to continue through the weekend.

Beam delivery update 11 December 2013

The cyclotron is delivering protons as needed to BL1B for proton irradiation and 9.8 µA to BL2A for RIB development with the new ion guide-laser ion source. PIF operation will continue until the end of the Schedule 125 running period while RIB development will continue through Friday.

Beam delivery update: Possible brief interruption(s) tomorrow 09 December 2013

The cyclotron is delivering a few thousand protons per second as needed to BL1B for PIF. BL2A will come back online tomorrow morning. There may be a brief interruption or two tomorrow morning for maintenance-type activities, but those will be coordinated with PIF and ISAC if necessary.

Beam delivery update 06 December 2013

The cyclotron is  delivering ~100 µA to BL1A and nA as needed to BL2C1 for PIF. BL1A will go offline for the year at 08:30 Saturday morning and beam will be switched to BL1B for the remainder of Schedule 125. BL2A will remain off over the weekend.

Beam delivery update 03 December 2013

Beam was injected around 17:30 this afternoon and the cyclotron is currently delivering ~95 µA to BL1A. Delivery to BL1A and to PIF on BL2C1 (on an as-needed basis) is scheduled to continue until Saturday morning.

Beam delivery update: No RIB, and a long maintenance tomorrow 02 December 2013

The cyclotron is currently delivering ~100 µA to BL1A and ~95 µA to BL2C4. This should continue until the start of maintenance tomorrow morning at ~07:00. BL2A is offline this week for ISAC target hall work.

Beam delivery update 28 November 2013

The cyclotron is running relatively well, delivering ~100 µA to BL1A, 60 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue through the weekend.

Beam delivery update 27 November 2013

The cyclotron is currently delivering ~100 µA to BL1A, 60 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue at least overnight.

Beam delivery update: Maintenance tomorrow 26 November 2013

The cyclotron is currently delivering ~100 µA to BL1A, 55 µA to BL2A, and ~90 µA to BL2C4. We should continue like this until the start of maintenance and ISIS development at 09:00 tomorrow.

Beam delivery update: No maintenance Tuesday 25 November 2013

Today’s efforts to find a controls workaround to the failure of the T2 left protect monitor plate appear to have been successful. The cyclotron is currently running and delivering ~100 µA to BL1A, 55 µA to BL2A, and ~95 µA to BL2C4.

Beam schedule change and T2 status 25 November 2013

BL1A was down for the weekend due to a target failure at T2. At the request of the µSR community, the “maintenance if required” shift scheduled for tomorrow has been postponed until Wednesday, November 27.

Beam delivery update: network disruptions this afternoon; recovery in progress 21 November 2013

Network problems starting around 15:30 disrupted beam delivery to various experiments and an apparent network glitch of some sort tripped a number of cyclotron subsystems at about 17:50. BL1A has been defined off due to vacuum issues that are still being sorted out while the beam currents in BL2A and BL2C4 have just been ramped back up to 55 and ~95 µA, respectively. UPDATE, 20:21: The cyclotron tripped off again at about 20:06 due to what appears to be a similar fault. There is no estimate at this time of when beam will be restored.

Beam delivery update 20 November 2013

The cyclotron is currently delivering ~110 µA to BL1A, 55 µA to BL2A, and ~95 µA to BL2C4; this is expected to continue through the weekend.

Beam delivery update 19 November 2013

ISIS beam development is still underway. That is expected to continue until ~20:30 as scheduled. Cyclotron areas have been locked and secured in preparation for injection this evening.

Beam delivery update: Beam development Tuesday 18 November 2013

The cyclotron’s delivering ~115 µA to BL1A, 55 µA to BL2A, and ~95 µA to BL2C4. BL2A is running at a reduced current in response to the needs of the current ISAC RIB experiment. Please note that there is beam development scheduled for tomorrow’s day shift (08:30 – 20:30).

Beam delivery update: Beam development Tuesday 15 November 2013

The cyclotron is delivering ~105 µA to BL1A, 70 µA to BL2A, and ~92 µA to BL2C4. This is expected to continue, with perhaps slightly higher currents in BL1A and 2C4, until Tuesday morning. Tuesday’s scheduled maintenance-if-required has been rescheduled as beam development; this is expected to take most, if not all, of the twelve-hour shift.

Beam delivery update 13 November 2013

The cyclotron is running normally and delivering ~107 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until next week.

Beam delivery update 12 November 2013

Today’s ISIS backplate change was complete by about 16:30 this afternoon and the cyclotron is now delivering ~105 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is scheduled to continue for the remainder of the week.

Beam delivery update (SPECIAL LONG WEEKEND EDITION) 08 November 2013

The cyclotron is currently delivering ~110 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until Monday, at which point PIF operation will resume on BL2C1.

Beam delivery update 07 November 2013

The cyclotron is running normally and delivering ~130 µA to BL1A, 70 µA to BL2A, and nA as needed to BL2C1 for PIF. This is still expected to continue through the weekend.

Beam delivery update 06 November 2013

The cyclotron is currently delivering ~130 µA to BL1A and 70 µA to BL2A. BL2C1 is off for the night. Delivery will continue with high currents in BL1A and BL2A and nA as needed for PIF use through the weekend. There will be no access to the ISAC-II experimental hall for ~4 hours starting at 09:00 tomorrow (Thursday). 

Beam delivery update 05 November 2013

The cyclotron has been delivering ~110 µA to BL1A, 70 µA to BL2A, and ~90 µA to BL2C4 since about 17:30. Some tuning is still needed, so the BL1A and BL2C4 currents may change somewhat overnight. 

Beam delivery update: Maintenance tomorrow 04 November 2013

The cyclotron is currently delivering ~110 µA to BL1A, 70 µA to BL2A, and ~95µA to BL2C4. This is expected to continue overnight. Beam will go off or maintenance at ~08:30 for ~8 hours.

Beam delivery update: Weekend edition 01 November 2013

The cyclotron is currently delivering ~275 µA to three beamlines: ~130 µA to BL1A, 50 to BL2A, and ~95 to BL2C4. BL2A is expected to be stepped up to 65 µA overnight (with a corresponding drop in the BL1A current).

Beam delivery update: potentially long maintenance tomorrow 30 October 2013

The cyclotron’s delivering ~130 µA to BL1A, 14 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until 08:30 tomorrow morning. At that time, the cyclotron will be shut off while work is done to repair a broken permanent magnet assembly in the ISIS I1 H- source.

Beam delivery update 29 October 2013

The cyclotron is running normally and delivering ~115 µA to BL1A, 14 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until the start of Thursday’s planned ISIS maintenance.

Beam delivery update: schedule changes later this week 28 October 2013

The cyclotron went off shortly after noon today because of a water leak in the main magnet power supply. Beam has been injected and Ops are in the process of tuning it through the machine. The plan is to restore high-current delivery — ~125 µA in BL1A, 14 µA in BL2A, and ~95 µA in BL2C4 — as promptly as possible.

Beam delivery update 24 October 2013

The cyclotron’s running normally, delivering ~130 µA to BL1A, 14 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue overnight.

Beam delivery update 22 October 2013

The cyclotron is currently delivering ~125 µA to BL1A, 14 µA to BL2A, and ~97 µA to BL2C4. This is expected to continue through tomorrow and into Thursday.

Beam delivery update 21 October 2013

The cyclotron is currently delivering ~125 µA to BL1A, ~90 µA to BL2C4, and 14 µA to BL2A; at ISAC, 10C at 6 MeV/u is being delivered to IRIS for experiment S1396. Maintenance is scheduled to begin at 08:30 tomorrow (Tuesday) and is expected to take 6–8 hours.

Beam delivery update 18 October 2013

For the first time in a while, there are actually three beamlines in operation: BL1A is receiving ~125 µA; BL2A, 14 µA; and BL2C4, ~95 µA. This is expected to continue through the weekend and until Tuesday's scheduled maintenance.

Beam delivery update 17 October 2013

The cyclotron is currently delivering ~130 µA to BL1A and ~95 µA to BL2C4. The BL1A current is somewhat higher than it's been in anticipation of BL2A coming online. That should happen shortly.

Beam delivery update 16 October 2013

BL1A is taking ~125 µA and BL2C4, ~95 µA. This is expected to continue overnight. ISAC and BL2A are expected to be ready for proton beam some time tomorrow.

Beam delivery update 15 October 2013

The current in BL1A has been increased to ~120 µA at the request of M15 experimenters; the cyclotron is also delivering ~95 µA to BL2C4 for isotope production. This is expected to continue overnight and tomorrow. BL2A remains off due to ongoing problems at ISAC. 

Beam delivery update (SPECIAL LONG WEEKEND EDITION) 11 October 2013

As has been the case for the last few days, the cyclotron is delivering about 110 µA to BL1A and nA as needed to BL2C1 for PIF. That is scheduled to continue until Tuesday morning. BL2A will remain off until at least Tuesday as efforts continue to repair the ISAC target hall crane.

Beam delivery update 10 October 2013

The cyclotron is still delivering ~110 µA to BL1A and nA as needed to BL2C1 for PIF. BL2A remains off pending further investigation of the west target station heat shield cooling and repairs to the ISAC target hall crane.

Beam delivery update 09 October 2013

The cyclotron is currently delivering ~110 µA to BL1A and nA as needed to BL2C1; that is scheduled to continue through the weekend. BL2A remains off due to problems with the ISAC target station.

Beam delivery update 08 October 2013

The cyclotron is currently delivering ~115 µA to BL1A and ~95 µA to BL2C4.  BL2A is off due to the apparent failure of a paddlewheel-type water flowmeter in the ISAC west target station and will remain off until the meter can be replaced.

Beam delivery update 01 October 2013

Maintenance wrapped up around 17:30 this afternoon and the cyclotron is delivering ~110µA to BL1A, 9.8 µA to BL2A, and ~60 µA to BL2C4. The current in BL2C4 is presently limited by the resistivity of the water package but will likely be increased as the night progresses.

Beam delivery update 30 September 2013

The cyclotron is currently delivering ~110 µA to BL1A and ~98 µA to BL2C4. BL2A is offline until approximately midnight as there is no need for RIB until then. Finally, the cyclotron is expected to be off for ~8–10 hours starting at 08:30 tomorrow to accommodate an ISIS backplate change.

Beam delivery update 23 September 2013

The cyclotron is currently running and delivering ~110 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight.

Beam delivery update 17 September 2013

The cyclotron is running normally and delivering ~120 µA to BL1A and ~95 µA to BL2C4. BL2A is offline; 8 µA will be requested shortly after tonight's 23:00 shift change.

Beam delivery update 13 September 2013

The cyclotron is currently delivering ~125 μA to BL1A and ~95 μA to BL2C4. BL2A was at 4 μA earlier today before the failure of a water-flow interlock forced it offline; that is being resolved and we expect to begin restoring the ISAC target heaters shortly in preparation for resuming BL2A operation sometime tonight.

Beam delivery update 03 September 2013

Today is the last day of beam delivery in Schedule 124. BL1A was taken offline earlier today and the cyclotron is now being used for BL1B development while still delivering ~45 µA to BL2C4 for isotope production.

Beam delivery update 29 August 2013

The cyclotron is currently delivering ~110 µA to BL1A and ~95 µA to BL2C4; that is expected to continue through the upcoming long weekend.

Beam delivery update 27 August 2013

The cyclotron is currently delivering ~130 µA to BL1A and ~95 µA to BL2C4 for isotope production; this will continue overnight then resume late tomorrow evening or early Thursday morning following cyclotron beam development.

Beam delivery update: BL1A operation extended through Labour Day 22 August 2013

Next week's PIF operation on BL1B has been reduced to two shifts. In the meantime, cyclotron operation is continuing as planned with ~130 µA being delivered to BL1A and ~95 µA to BL2C4.

Beam delivery update 20 August 2013

The cyclotron was off for about one hour this morning for scheduled maintenance; it's currently delivering ~130 µA to BL1A and ~90 µA to BL2C4. This is expected to continue through tomorrow.

Beam delivery update 19 August 2013

RIB delivery at ISAC has been cancelled for the remainder of Schedule 124. The cyclotron, meanwhile, continues to operate. It's delivering ~130 µA to BL1A and ~95 µA to BL2C4. A brief (one-hour) maintenance is expected to start at ~08:30 tomorrow.

Beam delivery update 22 July 2013

The cyclotron is currently delivering ~130 µA to BL1A, 8 µA to BL2A, and ~30 µA to BL2C4. This is expected to continue overnight…

Beam delivery update 17 July 2013

The cyclotron is delivering ~130 µA to BL1A, 6 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight, as is ISAC RIB delivery of 98Rb/Sr to the 8π for experiment S1376…

Beam delivery update 16 July 2013

Today's scheduled maintenance has been completed and the cyclotron is currently delivering ~125 µA to BL1A and ~98 µA to BL2C4. BL2A remains off as ISAC recovers from the failure of a cooling water pump. That's only expected to take another hour or so, at which point BL2A should resume operation at 6 µA…

Beam delivery update 15 July 2013

The cyclotron is currently running and delivering ~120 µA to BL1A, 6 µA to BL2A, and ~95 µA to BL2C4. Another I1 filament change is planned for tomorrow, along with a continued effort to find the cause of the last few filaments' rapid degradation…

Beam delivery update 12 July 2013

The cyclotron was down 15:00 this afternoon until about 19:30 following a power bump. The plan is to deliver 110+ µA to BL1A, 8 µA to BL2A, and ~95 µA to BL2C4 over the weekend…

Beam delivery update 11 July 2013

The cyclotron is currently delivering ~115 µA to BL1A, 8 µA to BL2A, and ~95 µA to BL2C4. At ISAC, yield measurements with the new UCx target are underway…

Beam delivery update 10 July 2013

Beam is slowly being restored. Full delivery to BL1A and BL2C4 is expected to resume within the next couple of hours…

Beam delivery update 09 July 2013

Most of today's planned maintenance has been completed, both on the cyclotron and at ISAC, but the stable sources group is still working on ISIS…

Beam delivery update 08 July 2013

The cyclotron is running and delivering ~110 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. Beam will go off at about 08:30 tomorrow for an ISIS backplate change…

Beam delivery update: weekend plans 05 July 2013

The cyclotron is currently delivering ~105 µA to BL1A, 50 µA to BL2A, and ~90 µA to BL2C4. This is expected to continue over the weekend through Monday…

Beam delivery update 04 July 2013

The ISIS I1 (ion source #1) filament appears to be failing despite having been replaced on Tuesday and the subsequent effort to identify possible leaks in the source vacuum system. Ops has adjusted split ratios to give 90~95 µA to BL1A, 50 µA to BL2A, and ~60 µA to BL2C4 in order to limit the impact on any one user or group of users. The plan is to run like this overnight, if possible…

Beam delivery update 02 July 2013

Maintenance is still underway as the ion source group continues to investigate a possible vacuum leak. Startup is expected around 20:30. At ISAC, a plan is in place to make use of the available RIB while recovery from this weekend's cryogenics failure continues…

ISAC cryogenics update 30 June 2013

The cryogenics group was able to cool all of the ISAC-II cryomodules using the Phase I system last night. The Phase II system, however, is still highly contaminated and Praxair has been contacted in an effort to obtain a fresh supply.

Beam delivery update 29 June 2013

The cyclotron's running smoothly, delivering 70 µA to BL2C4 for isotope production and ~25 µA to BL2A. BL1B operation has wrapped up for the day. There is no RIB user at ISAC due to a failure of the ISAC-II cryogenics system…

Beam delivery update [LONG WEEKEND EDITION] 28 June 2013

High-current operation with BL2A and BL2C4, coupled with low-current delivery to BL1B as needed, is scheduled to continue over the weekend. Unfortunately, the H- ion source filament that was installed on Tuesday appears to be dying and will be replaced this Tuesday, July 2…

Beam delivery update 27 June 2013

The cyclotron is running with slightly lower extracted currents than we had earlier in the week, ~110 µA in BL1A, 50 µA in BL2A, and ~95 µA in BL2C4. This is expected to continue until ~08:30 tomorrow morning.

Beam delivery update 25 June 2013

The cyclotron has been delivering ~130 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4 for about the last two hours. This is expected to continue until Friday morning.

Beam delivery update 24 June 2013

We've been delivering ~130 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4 since Friday night with only minor interruptions and will continue to do so until the start of maintenance at ~08:30 tomorrow. Tomorrow's maintenance is expected to be a full 6–8 hours to accommodate an ISIS backplate change…

Beam delivery update 21 June 2013

The cyclotron is running normally and delivering ~130 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue until the start of Tuesday's scheduled maintenance.

Beam delivery update 19 June 2013

The cyclotron's currently delivering ~130 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This will continue overnight and tomorrow…

Beam delivery update 17 June 2013

After a somewhat eventful day, the cyclotron is delivering ~125 µA to BL1A and 50 µA to BL2A…

BL2A vacuum failure RESOLVED 17 June 2013

The problem was found not to be a leak but rather a faulty controller for the last turbo pump on BL2A…

BL2A vacuum failure 17 June 2013

The cyclotron tripped off at about 02:30 this morning due to a vacuum failure in BL2A…

Beam delivery update 14 June 2013

The cyclotron is currently delivering ~130 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue, with occasional brief interruptions for access to the ISAC mass separator room, until the start of BL2C1 PIF operation on Monday…

Beam delivery update 13 June 2013

ISAC resumed operation shortly after 21:00 following last night's power supply failure. The cyclotron has been delivering ~130 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4 since…

Beam delivery update 12 June 2013

RIB delivery at ISAC was interrupted earlier this evening by the failure of the target heater power supply. BL2A is off as a result. The cyclotron is still running, however, and delivering ~135 µA to BL1A and ~95 µA to BL2C4…

Beam delivery update 11 June 2013

Maintenance day has been finished since late afternoon and the cyclotron is delivering ~130 µA to BL1A, 70 to BL2A, and ~97 to BL2C4…

Beam delivery update 10 June 2013

The cyclotron is delivering ~130 µA to BL1A, 70 µA to BL2A, and ~95 µA to BL2C4. This will continue until 08:30 tomorrow morning, at which point BL2A will go off for ~4 hours for ISAC maintenance…

Beam delivery update 07 June 2013

The cyclotron is currently delivering ~130 µA to BL1A, 70 to BL2A, and ~95 to BL2C4. This is expected to continue over the weekend and through Monday…

Beam delivery update [CORRECTED] 06 June 2013

The cyclotron is currently delivering ~260 µA total, ~115 to BL1A, 50 to BL2A, and ~95 to BL2C4. The BL2A current will be stepped up to 70 µA overnight as the ISAC target vacuum allows…

Beam delivery update 04 June 2013

Today's scheduled maintenance is well underway with many of the major jobs (tank defrost, ISIS backplate change, electrical work in the proton hall, etc.) having already been completed…

Beam delivery update 03 June 2013

The cyclotron is currently delivering ~90 µA to BL1A, 45 µA to BL2A, and ~95 µA to BL2C4. Delivery to all three high-current beamlines will continue until the start of a 36-hour maintenance period at 08:30 tomorrow morning…

Beam delivery update 31 May 2013

BL2A has been defined off to allow access to the mass separator room while the ISAC Controls group works to fix a controls communication issue. In the meantime, the cyclotron is delivering ~110 µA to BL1A and ~98 µA to BL2C4…

Beam delivery update 29 May 2013

The cyclotron is currently delivering 45 µA to BL2A and ~95µA to BL2C4…

Maintenance tomorrow 27 May 2013

Maintenance is scheduled to begin ~08:30 tomorrow (Tuesday) morning…

Beam delivery update 24 May 2013

The cyclotron is currently running and delivering ~105 µA to BL1A and 45 µA to BL2A while BL2C is set up for proton therapy on BL2C1. That will continue over the weekend, starting at 07:00 tomorrow (Saturday), 07:30 Sunday, and the usual time of 08:00 Monday…

Beam delivery update 23 May 2013

Very little has changed in the last 24 hours: the cyclotron is delivering ~100 µA to BL1A, 45 µA to BL2A, and ~90 µA to BL2C4, while at ISAC 11Be is being delivered to TIGRESS for experiment S1297…

Beam delivery update 22 May 2013

Cyclotron operation resumed yesterday evening and the machine is currently delivering ~95 µA to each of BL1A and BL2C4 and 45 µA to BL2A…

Beam delivery update 21 May 2013

ISIS beamline tuning is underway following this morning's backplate change. The rest of today's scheduled maintenance has been completed and we're just waiting for beam to be injected at this point…

Maintenance tomorrow 20 May 2013

There are 8–12 hours of maintenance scheduled for tomorrow (Tuesday) to accommodate an ISIS backplate change. Beam is expected to go off at ~08:30…

Beam delivery update (SPECIAL LONG WEEKEND EDITION) 17 May 2013

Beam ran relatively well to BL1A, BL2A, and BL2C4 overnight. PIF operation on BL2C1 resumed at about 10:30 this morning, but about half an hour later the cyclotron tripped off due to a controls problem…

Beam delivery update 16 May 2013

The cyclotron is currently delivering ~100 µA to BL1A, 45 µA to BL2A, and ~95 µA to BL2C4. 11Be is being delivered to TIGRESS for experiment S1297 but the yield is significantly lower than expected…

Beam delivery update 15 May 2013

Maintenance was completed early this afternoon and the cyclotron is currently delivering ~110 µA to BL1A, 45 µA to BL2A, and nA as needed to BL2C1 for PIF…

Beam delivery update 13 May 2013

The cyclotron is currently delivering ~100 µA to BL1A, 45 µA to BL2A, and ~95 µA to BL2C4 for isotope production…

Beam delivery update 09 May 2013

The cyclotron has actually run rather well since early this morning, delivering >100 µA to BL1A, 40–45 µA to BL2A, and 95–100 µA to BL2C4 for much of the day…

Beam delivery update 08 May 2013

It's been a rough day. Ops haven't been able to get more than about 190 µA total from the cyclotron into BL1A, 2A, and 2C4…

Beam delivery update 07 May 2013

Maintenance was completed earlier today and the cyclotron was delivering ~50 µA to BL1A and ~75 µA BL2C4 for much of the afternoon…

Beam delivery update 06 May 2013

The cyclotron is delivering 5 µA to BL2A for ISAC target scans and ~94 µA to BL2C4 for isotope production. Beam development is expected to start around 20:00 this evening and continue through the night…

Beam delivery update 03 May 2013

The cyclotron is running and delivering 40 µA to BL2A and ~70 to BL2C4. Also: road closures Sunday…

Beam delivery update 02 May 2013

After a bit of a slow startup following yesterday's RF repairs, the cyclotron is running and delivering 40 µA to BL2A and ~95 µA to BL2C4…

Beam delivery update 01 May 2013

After running overnight with beam in BL2A, BL2C4, and BL1A (albeit at low currents) the cyclotron tripped off at 09:34 this morning due to an RF failure…

Beam delivery update 30 April 2013

Maintenance day is largely over and the cyclotron is starting back up…

Beam delivery update 29 April 2013

The cyclotron is currently delivering 40 µA to BL2A and ~90 µA to BL2C4 for isotope production at STF…

Beam delivery update 26 April 2013

The cyclotron is currently delivering ~40 µA to BL2A…

Beam delivery update 23 April 2013

Today's planned cyclotron development was derailed by a water leak in the main magnet power supply. That has been fixed and the cyclotron is gradually being restored…

Beam delivery update 22 April 2013

The cyclotron is running – BL2A saw ~4.5 µA for most of the weekend – but is still very much in startup mode…

Beam delivery update 19 April 2013

The cyclotron RF is currently operating at 86 kV and beam has been injected…

Beam delivery update 18 April 2013

Repairs to the cyclotron system will continue tomorrow. Some testing with HV was carried out today but more tuning is required…

Beam delivery update 15 April 2013

Many of the repairs to the cyclotron RF system have been completed but the 480 V contactor still has to be replaced…

Beam delivery update 12 April 2013

Repairs to the cyclotron RF system are still expected to be completed next week…

Beam delivery update 10 April 2013

Stable-beam delivery is continuing at ISAC while repairs to the cyclotron RF system are still expected to carry on into next week…

Beam delivery update 08 April 2013

Repairs to the cyclotron RF system in the wake of Friday's failure are expected to delay startup of the cyclotron RF to about April 19…

Beam delivery update 05 April 2013

One of the cyclotron RF intermediate power amplifiers sparked over this afternoon, destroying a 480 V breaker and damaging components inside the amplifier cabinet…

Beam delivery update: We're back. 04 April 2013

18O5+ from the OLIS multi-charge ion source (i.e. the Supernanogan) was tuned through the TIP chamber in ISAC-I earlier today…

Beam delivery update (CORRECTED) 02 April 2013

Shutdown activities are drawing to a close. The cyclotron is under vacuum, the main magnet is on, and after a delay over the weekend the RF has been started and is slowly being conditioned…

Shutdown 2013 07 January 2013

The lab is now into its annual winter shutdown…

End-of-year update 21 December 2012

It is very quiet at the lab right now…

Beam delivery update 20 December 2012

Thanks to the efforts of a large number of people in pretty much every support group, the cyclotron is currently delivering 50 µA to BL2A2 and nA as needed to BL2C1. Shutdown effectively starts at the close of business tomorrow…

Update 20 December 2012

Recovery is proceeding more quickly than expected. It may be possible to inject beam into the cyclotron during the day Thursday…

Winter weather 19 December 2012

File this under things you don't often see at TRIUMF…

Beam delivery update 19 December 2012

We are just beginning to recover from a site-wide power outage…

Beam delivery update 17 December 2012

Current conditions: 50 µA to BL2A2 and nA as needed to BL1B or BL2C1 for PIF. Tomorrow's scheduled maintenance is expected to be brief, if it even happens…

Beam delivery update 14 December 2012

Current conditions: 50 µA to BL2A, with BL1B idle for the night. The final RIB experiment of the year is underway…

Beam delivery update 13 December 2012

Current conditions: 50 µA to BL2A2. PIF will resume on BL1B tomorrow…

Beam delivery update 12 December 2012

Current conditions: 5 µA to BL1A, 50 µA to BL2A2, and nA as needed to BL2C1. There is a brief maintenance period planned for tomorrow morning…

Beam delivery update 11 December 2012

Current conditions: 5 µA to BL1A, 50 µA to BL2A2, and nA as needed to BL2C1. BL1A will run at reduced currents until Thursday morning…

Beam delivery update 10 December 2012

Current conditions: 110 µA to BL1A, 50 µA to BL2A2, and nA as needed to BL2C1 for PIF. BL1A will cease high-current operation for the year at 08:30 tomorrow morning…

Beam delivery update 07 December 2012

Current conditions: 110 µA to BL1A, 50 µA to BL2A2, and ~75 µA to BL2C4. Power supply and vacuum failures prevented operation for several hours earlier today…

Beam delivery update 05 December 2012

Current conditions: 110 µA to BL1A, 50 µA to BL2A2, and nA as needed to BL2C1 for PIF. Final RIB setup to TUDA is underway…

Beam delivery update 04 December 2012

Current conditions: Scheduled maintenance, to be followed by beam development overnight…

Beam delivery update 03 December 2012

Current conditions: 110 µA to BL1A and 50 µA to BL2A2, with BL2C1 configured for proton therapy. PT, maintenance, and beam development tomorrow…

Correction 30 November 2012

Proton therapy will start at ~08:00 Saturday–Tuesday, not 08:30 as indicated yesterday. The Beam Delivery group regrets the error.

Beam delivery update 29 November 2012

Current conditions: 110 µA to BL1A, 50 µA to BL2A2, and ~70 µA to BL2C4. There is proton therapy scheduled for tomorrow and over the weekend…

Beam delivery update 28 November 2012

Current conditions: ~110 µA to BL1A, 50 µA to BL2A2, and ~75 µA to BL2C4, but things have been unstable for most of the day. Also: Last chance to support Movember…

Beam delivery update 27 November 2012

Current conditions: 115 µA to BL1A, 25 µA to BL2A2, and ~75 µA to BL2C4, with ISAC target startup underway…

Beam delivery update 26 November 2012

The cyclotron is currently delivering ~115 µA to BL1A and ~75 µA to BL2C4. BL2A has been off since about 02:00 this morning…

Beam delivery update 23 November 2012

The cyclotron is currently delivering 115 µA to BL1A, 9.8 µA to BL2A3, and ~75 µA to BL2C4, and something is causing the cyclotron active sump to fill at an increased rate…

Beam delivery update 21 November 2012

The cyclotron is still delivering 115 µA to BL1A, 9.8 µA to BL2A3, and ~75 µA to BL2C4…

Beam delivery update 20 November 2012

Everything is back to normal following today's scheduled maintenance…

Beam delivery update 19 November 2012

The cyclotron is currently delivering 115 µA to BL1A, 9.8 µA to BL2A3, and ~75 µA to BL2C4…

Beam delivery update 16 November 2012

The cyclotron is currently delivering 115 µA to BL1A, 6 µA to BL2A3, and ~75 µA to BL2C4. Also: a Movember update…

Beam delivery update 14 November 2012

Maintenance wrapped up this morning and the cyclotron has been running since early afternoon…

Beam delivery update 13 November 2012

Maintenance began this morning as scheduled and will continue overnight…

Beam delivery update 12 November 2012

ISAC-II operation is still suspended due to cryogenics problems…

Beam delivery update 11 November 2012

DSAM experiment S1378 at ISAC-II is currently on hold due to vacuum and cryogenics failures…

Beam delivery update 09 November 2012

The cyclotron is currently delivering ~115 µA to BL1A, 9.8 µA to BL2A3, and ~70 µA to BL2C4. We don't anticipate any significant changes prior to the start of maintenance on Tuesday…

Site air leak 06 November 2012

There appears to be a large leak somewhere in the site compressed air system…

Beam delivery update 06 November 2012

[UPDATED] Right now, we're delivering ~115 µA to BL1A, 9.8 µA to BL2A3, and ~78 µA to BL2C4…

Beam delivery update 05 November 2012

Things are running relatively smoothly with the cyclotron currently delivering ~115 µA to BL1A, 9.8 µA to BL2A3, and ~75 µA to BL2C4…

Beam delivery update 02 November 2012

The cyclotron is currently delivering ~120 µA to BL1A, 10 µA to BL2A3, and ~50 µA to BL2C4…

Beam delivery update 30 October 2012

Cyclotron beam development is underway and will continue overnight…

Beam delivery update 29 October 2012

The cyclotron's delivering ~115 µA to BL1A, 40 µA to BL2A2, and ~50 µA to BL2C4. Cyclotron beam development is planned for tomorrow…

Beam delivery update 26 October 2012

The cyclotron is currently delivering ~115 µA to BL1A and 40 µA to BL2A2. Also: Movember

Beam delivery update 25 October 2012

BL1A is currently running at ~100 µA, with BL2A2 at 40 µA and BL2C1 taking nA as needed

Beam delivery update 24 October 2012

The BL1A tunnel has been secured and is ready for beam…

BD updates by e-mail 24 October 2012

Beam delivery updates are also available by e-mail…

Beam delivery update 23 October 2012

Maintenance day is over and the cyclotron is currently delivering 40 µA to BL2A2 and nA as needed to BL2C1 for PIF…

Beam delivery update 18 October 2012

RIB development and yield measurements are underway at ISAC…

Beam delivery update 17 October 2012

Right now, the cyclotron is delivering 40 µA to BL2A2 and ~45 µA to BL2C4…

Subscribe to Group News